Provide answers for the following questions.

Provide answers for the following questions for each problem.

a) What is the EPQ?

b) What is the maximum inventory? What is the annual holding cost?

c) What is the number of runs or batches (setup) per year? What is the annual setup cost?

d) What is the total annual cost?

1) The Dine Corporation is a producer of brass couplings. The firm operates 220 days per year and uses (demand rate) the couplings at a steady rate of 50 per day (therefore making annual demand = 11,000 ie 220 x 50). Couplings can be produced at a rate of 200 per day. Annual storage cost is $2 per coupling, and the machine setup cost is $70 per run.
Answer #1

Economic Production Quantity is the number of unit that is added to the inventory which minimize the total inventory cost. It maintain a balance between ordering costs and carrying costs.

m = 220 days d=50/day D = 11000 P = 200/day H = $2/ coupling / year S = $70 oeun EPS = 2DS X told H | 2 x 11000 x 70 / 200 waTotal cost = Annual selup & Annual carrying costa cost 2 TC (oxs) + ( I mare x H) (1000 x 70) + ( 160 x 2) 760.11 + 760 $1520